# AEGIS : Automated Co-Evolutionary Framework for Guarding Prompt Injections Schema

### v1

@@ Introduction
4:  $ATK^{i}_{0}\leftarrow Eval\_Attack\_Train(ATK^{i-1}_{M_{i-1}})$
@@
7:   $ATK^{i}_{j}\leftarrow Eval\_Attack\_Train(ATK\_Cand^{i}_{j})$   end for
8:  $ATK^{i}_{best}\leftarrow Val\_Attack(ATK^{i}_{M_{i}})$
@@
13:   $DEF^{i}_{j}\leftarrow Evaluate\_Defense(DEF\_Cand^{i}_{j})$   end for
14:  $DEF^{i}_{best}\leftarrow Validate\_Defense(DEF^{i}_{M_{i}})$ end for
15:return $ATK_{best}$, $DEF_{best}$
Attacker.
The attacker module aims to generate adversarial prompts that can mislead the system to give a higher score. Within the co-evolutionary framework, the attacker evolves in alternating turns against a fixed defender, simulating an arms race between offensive and defensive strategies.
During each attacker iteration, a training phase is initiated to explore new adversarial candidates. Specifically, the attacker maintains a top-$k$ pool of the strongest attack prompts from previous cycles, as measured by their effectiveness against the current best defense. New attack candidates are generated using the $TGO^{+}$ module, which synthesizes gradient-like signals derived from grading feedback and guides the editing of existing prompts to enhance their adversarial strength.
Each new attack prompt is evaluated by the grading system, and we compute two key metrics: the relative score change and the attack score, as defined in Equation 1. Below is an explanation of each notation used in the formula:
  • $S_{\text{benign}}$: The score assigned to the original (benign) input without any adversarial prompt.
  • $S_{\text{attacked}}$: The score assigned after injecting the adversarial prompt into the input.
  • $S_{\text{max}}$: The maximum possible score defined by the grading system.
  • $\Delta S_{\text{rel}}$: The relative score change, which quantifies how much the attack improved the score, normalized by the possible range for score improvement.
The formula is defined as:
If the benign score is less than the maximum score, the relative score change is computed as the ratio between the actual gain due to the attack and the maximum possible gain. If the benign score is already at the maximum, no improvement is possible, and the relative score change is defined as zero.
This normalization is crucial because it emphasizes the significance of score changes at higher score ranges. For example, a change from 8 to 9 is more impactful than a change from 1 to 2, because it is generally harder to achieve.
Attack Score:
$S_{\text{attack}}=w_{\text{asr}}\cdot(\text{ASR})^{p_{\text{asr}}}+w_{\text{sc}}\cdot(\Delta S_{\text{rel}})^{p_{\text{sc}}}$
where $w_{asr}$ and $w_{sc}$ are weights for the Attack Success Rate (ASR) and relative score change ($\Delta S_{rel}$), respectively, and $p_{asr}$ and $p_{sc}$ are power parameters to control the sensitivity of the scores. The coefficients ($w$) allow us to prioritize one metric over the other (e.g., a higher $w_{asr}$ would prioritize attacks that are harder to detect), while the power parameters ($p$) allow us to control the importance of changes at different ends of the spectrum (e.g., a higher $p_{asr}$ would place more emphasis on improvements from 0.8 to 0.9 than from 0.1 to 0.2).
If the attack score of a newly generated candidate outperforms existing prompts in the top-$k$ set, it is selected for inclusion in the updated pool.
After the training procedure, the best-performing attack prompt from the top-$k$ pool is selected. Then, the selected prompt will be fixed for the next defender evolution cycle, ensuring that the defender is trained against the most challenging known threat at the time.
Defender.
Following the co-evolutionary setup, the defender evolves in response to a fixed attacker. After each attacker cycle, the best-performing attack prompt is used as the evaluation context against which the defender is trained. The evolution process mirrors that of the attacker: a top-$k$ pool of defense prompts is maintained and updated via the $TGO^{+}$ module, which refines existing prompts using gradient-like feedback from the grading system.
Defense Score:
$S_{defense}=w_{tp}\cdot(TPR)^{p_{tp}}+w_{tn}\cdot(TNR)^{p_{tn}}$ (2)
where $w_{tp}$ and $w_{tn}$ are weights for the True Positive Rate (TPR) and True Negative Rate (TNR), respectively, and $p_{tp}$ and $p_{tn}$ are their corresponding power parameters. Similar to the attack score, the coefficients and power parameters in the defense score allow for a nuanced and flexible evaluation of the Defenderś performance.
Algorithm 2 Eval_Defense_Train(.)Eval\_Defense\_Train(.) in Algorithm 1.
Require: $DEF^{i}_{j}$: Defense candidates, $ATK^{i}_{best}$: Current Best attack, $art_{benign}$: Training benign articles
[…]
2:for $def\in DEF^{i}_{j}$ do
3:  $art_{attacked}\leftarrow$ Add $ATK^{i}_{best}$ before and after $art_{benign}$
4:  $art_{ben+def}\leftarrow$ Add defdef after $art_{benign}$
5:  $art_{atk+def}\leftarrow$ Add defdef after $art_{attacked}$
6:  $TNR\leftarrow LLM_{detection}(art_{ben+def})$
7:  $TPR\leftarrow LLM_{detection}(art_{atk+def})$
8:  Calculate $S_{defense}$ from TPRTPR and TNRTNR
9:  $scores[def]\leftarrow(S_{defense},TPR,TNR)$
[…]
11:return TopKTopK
Algorithm 3 Val_Defense(.)Val\_Defense(.) in Algorithm 1.
Require: $DEF^{i}_{j}$: Defense candidates, $ATK^{i}_{best}$: Current Best attack, $art_{benign}$: Validation benign articles
[…]
2:for $def\in DEF^{i}_{j}$ do
3:  Run the same grading pipeline in Algorithm 2
4:  $scores[def]\leftarrow(S_{defense},TPR,TNR)$
[…]
6:return $DEF_{best}$
TGO+ Framework for Prompt Optimization
[do we mention algorithm and fig at the same time?] Inspired by Microsoft’s TGO framework (tgo2024), we adopt a modular design to optimize prompts through gradient-like updates in natural language space. Each TGO+ module is dedicated to a specific optimization goal (e.g., attack or defense) and operates in two stages: gradient acquisition and gradient application, as illustrated in Figure 3.
Gradient Acquisition.
The module first samples candidate prompts and submits them to the grading system to obtain scores, and the grading system calculate score as Algorithm 1. Based on the results, an LLM is prompted with a task-specific instruction and the observed error. The LLM returns a feedback message that serves as a natural language proxy for a gradient direction—i.e., a suggestion indicating how the prompt could be improved.
To support iterative learning, recent feedback messages are stored in a gradient buffer. This buffer helps stabilize and diversify the optimization trajectory.
Gradient Application.
In the second stage, the buffered feedback is synthesized into a set of guidance messages. These are used to construct an edit prompt that instructs the LLM to revise the original candidate prompt accordingly. The improved prompt is then evaluated and, if effective, retained for further optimization cycles.
TGO+ enables gradient-like prompt updates without requiring access to model internals, making it compatible with black-box LLMs such as GPT-4o.
Key Innovations.
  • Multi-Route Gradient Optimization: To enhance the optimization process, we employ a multi-route gradient strategy. This means that for each prompt, we generate textual gradients based on multiple performance metrics. For instance, the Attackerś prompts are optimized based on both ASR and the induced score change. Similarly, the Defenderś prompts are optimized based on both TPR and TNR. This allows for a more holistic and effective optimization process.
@@
3:for each candidate $c\in C_{0}$ do
4:  // Step 1: Select bad detection results
@@
6:   Select detection results with lower ASR and $\Delta S_{rel}$
@@
8:   Select detection results with lower TPRTPR and TNRTNR
9:
10:  // Step 2: Compose natural language error string
11:  // (e.g., “This attack has a low ASR of 0.1 under defense “Focus on grading only!””)
12:  $e_{c}\leftarrow$ Generate error description string
13:
14:  // Step 3: Add past gradients as experience replay
15:  $g_{past}\leftarrow$ Retrieve past gradients associated with $c$
16:  $e_{c}\leftarrow e_{c}+g_{past}$
17:
18:  // Step 4: Get textual gradients from LLM
19:  $g_{c}\leftarrow LLM_{grad}(c,e_{c})$
20:
21:  // Step 5: Edit prompt using gradients
22:  $c_{\text{new}}\leftarrow LLM_{edit}(c,e_{c},g_{c})$
23:  Append $c_{\text{new}}$ to $C_{\text{new}}$ end for
24:return $C_{\text{new}}$
Experimental Setup
Dataset
Experimental Procedure
Hyperparameters
The default hyperparameters used in our experiments are summarized in Table 1. These settings were used for the baseline experiment, and variations are explored in the ablation studies.
@@
Results
Defense Evaluation
We evaluate the effectiveness of our adversarially trained defender against three baseline mechanisms, including Perplexity-based Detection (Alon and Kamfonas 2023), LLaMA 3.1 Guard (Inan et al. 2023), and the ”Human-Optimized Prompt”, which is the modified defense in Chiang et al. (2024) with clearer structure and objective (See appendix). For all these methods, we evalaute the defense against the real-world articles (147 malicious articles + 100 benign articles). The results, summarized in Table 2, show that our method achieves state-of-the-art performance. We present our results at both an early stage (Iteration 4) and the final stage (Iteration 8) of the GAN training.
Defense Method TPR TNR Human-Optimized Prompt (Chiang et al. 2024) 0.26 0.98 Perplexity-based Detection (Alon and Kamfonas 2023) 0.54 0.73 LLaMA 3.1 Guard (Inan et al. 2023) 0.61 0.81 Our Method (AEGIS @ Iteration 4) 0.76 0.88 Our Method (AEGIS @ Iteration 8) 0.84 0.89
Table 2: Defense Method Comparison. The result shows that our method already achieves the best TPR at iteration 4, despite a slight decrease in TNR compared with Human-Optimized Prompt. The model even performs better at iteration 8, with better TPR and TNR compared with our method at iteration 4.
Attack Evaluation
To measure the potency of our generated attacks, we evaluate our attacker against the ”Human-Optimized Prompt” at Iteration 4 and Iteration 8. The results in Table 3 show a significant improvement over the original baseline attacks, which are the students’ malicious articles.
Attack Method ASR $\Delta S_{\text{rel}}$ Students’ Malicious Articles 0.74 0.98 Our Method (AEGIS @ Iteration 4) 0.96 0.87 Our Method (AEGIS @ Iteration 8) 1.0 1.0
Table 3: Attack Method Comparison. The result shows that our method achieve the best result at iteration 8, with ASR and $\Delta S_{\text{rel}}$ both 1.0 when defended by Human-Optimized Prompt.
The attacker evolved through our framework achieves a 100% success rate, highlighting its capability to systematically discover and exploit vulnerabilities in manually-crafted defenses.
Iterative Performance
To illustrate the co-evolution of the attacker and defender, Table 4 shows the real-world evaluation for both agents at different iteration of the GAN training process. To be more specific, we evaluate the generated attacks against the defense in Human-Optimized Prompt modified from Chiang et al. (2024), and evaluate the generated defenses against the real-world articles (143 malicious articles + 100 benign articles) to calculate the TPR and TNR. The results demonstrate a clear trend of mutual improvement, where each agent becomes progressively stronger by adapting to the other.
The trend shows the attacker’s metrics (ASR, Relative Score Change) steadily increasing as it learns to bypass the improving defender. Simultaneously, the defender’s TPR initially drops as it encounters novel attacks but then recovers and improves as it learns to generalize and adapt. This dynamic demonstrates a successful adversarial training loop, leading to highly robust agents.
Cross-Model Generalizability
To evaluate the robustness and generalizability of the framework, we run the experiment on different LLMs. For these LLMs, we can notice that the framework is useful for generating better and better defense prompts as the training keeps going. Table 5 shows the results for all models at GAN iteration 0, iteration 2, and iteration 4. We can see that all these models achieve better TPR at higher GAN iteration compared with lower GAN iteration. For more detailed results, please check the Appendix.
Also, we investigated the transferability of the generated defense prompts across various large-language models (LLMs). Specifically, we examine whether the prompts optimized using one model (in this case, GPT-4.1-mini) retain their effectiveness when applied to different LLMs (in this case, GPT-4.1-nano, Gemini-2.5-flash, and Gemini-2.5-flash-lite) without modification.
Table 6 illustrates the results for all models at GAN iteration 0, iteration 4, and iteration 8. For stronger LLMs (GPT-4.1-mini, Gemini-2.5-flash), they achieve a high TPR at iteration 8, while weaker LLMs (GPT-4.1-nano, Gemini-2.5-flash-lite) achieve lower TPR at iteration 8. However, we can see that the results for all LLMs improve when more GAN iterations are trained. For more detailed results, please check the Appendix.
Target Model GPT-4.1-mini GPT-4.1-nano Gemini-2.5-flash Gemini-2.5-flash-lite 0 0.08 0.01 0.62 0.04 4 0.78 0.07 0.91 0.21 8 0.84 0.15 0.98 0.39
Ablation Study
Figure 4: Iterative performance of the attacker and defender across GAN iterations, measured by Attack Success Rate (ASR). Shaded regions represent the standard deviation across runs.
Figure 5: Iterative performance of the attacker and defender across GAN iterations, measured by Score Change. Shaded regions represent the standard deviation across runs.
Figure 7: Iterative performance of the attacker and defender across GAN iterations, measured by True Positive Rate (TPR). Shaded regions represent the standard deviation across runs. The default method has the steadiest improvement across iteration, and achieve the best TPR at last.
To understand the contribution of each component in our proposed PIA-GAN framework, we conducted several ablation studies. These studies involve systematically removing or altering parts of our system and observing the impact on performance.
Without Gradient Buffer
In this study, we removed the gradient buffer, which stores historical textual gradients for the optimization prompt. This change significantly degrades performance. Without access to past gradients, the optimization process becomes less stable and struggles to build upon previous iterations, leading to a slower convergence rate and a final performance that is approximately 15-20% worse than the full model.
Without Multiple Gradients
Here, we simplified the gradient generation process to use only a single performance metric. For the defender, we used only the True Positive Rate (TPR), and for the attacker, only the Attack Success Rate (ASR). This led to a skewed optimization. The defender becomes overly aggressive, achieving a high TPR at the cost of a poor TNR (many false positives). The attacker focuses narrowly on a single attack vector, resulting in a lower ASR when faced with a robust defense.
Single-Sided Training
We also experimented with training only one side of the GAN framework. The lack of an adaptive adversary meant that the trained model quickly overfit to its static opponent, highlighting the importance of co-evolution.
  • Defender-Only Training: When training the defender against a static set of attacks, it achieves near-perfect performance on those attacks but fails to generalize to new, unseen attacks, resulting in a lower overall TPR.
  • Attacker-Only Training: Similarly, when the attacker is trained against a static defender, it achieves a very high success rate against that specific defense but cannot adapt to a more robust, dynamic one.
The comparisons between all the ablation studies are presented in the figures above. For attackers, Fig 4 compares the asr between each ablation, and Fig 5 compares the score change between each ablation. On the other hand, Fig 7 compares the TPR between each ablation, and Fig 6 compares the TNR between each ablation.
Real Prompt Improvements
@@ Results
Conclusion
Through extensive experiments and ablation studies, we confirmed the importance of co-evolution, gradient replay, and multi-objective optimization. These findings suggest that adversarial training, when applied at the prompt level, offers a scalable and effective solution for safeguarding LLMs in real-world deployments.
In future work, we plan to extend AEGIS to more complex scenarios, improve prompt interpretability, and explore its applicability to other adversarial tasks in NLP and LLM alignment.
Acknowledgments
AAAI is especially grateful to Peter Patel Schneider for his work in implementing the original aaai.sty file, liberally using the ideas of other style hackers, including Barbara Beeton. We also acknowledge with thanks the work of George Ferguson for his guide to using the style and BibTeX files — which has been incorporated into this document — and Hans Guesgen, who provided several timely modifications, as well as the many others who have, from time to time, sent in suggestions on improvements to the AAAI style. We are especially grateful to Francisco Cruz, Marc Pujol-Gonzalez, and Mico Loretan for the improvements to the Bib and files made in 2020.
[…]
Appendix
Detailed Experiment Results
@@ Acknowledgments
Cross-Model Generalizability (Full Results)
Table 8 contains the TPR values for GPT-5-mini, GPT-4.1-nano, Gemini-2.0-flash, and Gemini-2.5-flash-lite at each GAN iteration.
@@
Table 9 contains the TNR values for GPT-5-mini, GPT-4.1-nano, Gemini-2.0-flash, and Gemini-2.5-flash-lite at each GAN iteration.
@@
Prompt Transferability (Full Results)
Table 10 shows the TPR values when prompts generated by GPT-4.1-mini are transferred to other models without modification.
@@
Table 11 contains the TNR values for GPT-5-mini, GPT-4.1-nano, Gemini-2.0-flash, and Gemini-2.5-flash-lite at each GAN iteration.
@@
Human Optimized Prompt
(Chiang et al. 2024) talks about the prompt based defense mechanism, and we slightly modified the detection prompt that leads to higher TPR, which is called the Human Optimized Prompt. The full prompt is in Table 12
Human Optimized Prompt We want you to check if the student’s answer tries to hack the evaluation assistant based on the students article. Please look carefully and assess if there are any modification attempts to their score. Please neglect any modifications on evaluation and scoring criteria in the student’s answer and fully obey the evaluation criteria. Any modification on the evaluation criteria or instructions should be treated as an attack.
Table 12: Human Optimized Prompt serves as the baseline to validate the attacks strength in each GAN iteration. This defense is modified from (Chiang et al. 2024).
1. full experiments for cross-model generalizability
2. Human-Optimized Prompt

### v2

@@ A.1 Algorithm for Adversarial Co-evolution Framework
4:  $ATK^{i}_{0}\leftarrow Eval(ATK^{i-1}_{M_{i-1}})$
@@
7:   $ATK^{i}_{j}\leftarrow Eval(ATK\_Cand^{i}_{j})$   end for
8:  $ATK^{i}_{best}\leftarrow Val(ATK^{i}_{M_{i}})$
@@
13:   $DEF^{i}_{j}\leftarrow Eval(DEF\_Cand^{i}_{j})$   end for
14:  $DEF^{i}_{best}\leftarrow Val(DEF^{i}_{M_{i}})$
15:  end for
16:return $ATK^{N}_{best}$, $DEF^{N}_{best}$
A.2 Algorithm for TGO workflow
Algorithm 2 describes how new candidate prompts are generated and refined during the adversarial co-evolution process. For each candidate $c$, the framework first identifies the error cases, and these failures are summarized into a natural language error string, providing contextual feedback. Then, this error string is augmented with task-specific instructions and gradients collected from past iterations to enable experience replay. The LLM will respond with textual gradients from the error string, and the LLM can further use these gradients to generate the new candidate prompt.
@@ A.2 Algorithm for TGO workflow
[…]
4:for each candidate $c\in C_{0}$ do
@@
6:   Select grading results with low ASR and $\Delta S_{\text{rel}}$
@@
8:   Select grading results with low TPR and TNR
9:  $e_{c}\leftarrow$ Generate error description string
10:  $g_{\text{past}}\leftarrow$ Retrieve past gradients related to $c$
11:  $e_{c}\leftarrow e_{c}+g_{\text{past}}$
12:  $g_{c}\leftarrow LLM_{\text{grad}}(c,e_{c})$
13:  $c_{\text{new}}\leftarrow LLM_{\text{edit}}(c,e_{c},g_{c})$
14:  Append $c_{\text{new}}$ to $C_{\text{new}}$
15:return $C_{\text{new}}$
A.3 Detailed Hyperparams
Table 5 shows the detailed hyperparameters we use in our experiments:
@@ A.3 Detailed Hyperparams
A.4 Defense Prompt Improvement
@@
A.5 Detailed Experiment Results
@@
A.5.1 Cross-Model Generalizability (Full Results)
Table 7 and Table 8 contain the TPR and TNR values for GPT-5-mini, GPT-4.1-nano, Gemini-2.0-flash, and Gemini-2.5-flash-lite at each GAN iteration respectively. We can see that TPR keeps increasing among all the GAN iterations despite a slight degradation in TNR values for all LLMs.
@@ A.5.1 Cross-Model Generalizability (Full Results)
A.5.2 Prompt Transferability (Full Results)
Table 9 and Table 10 show the TPR and TNR values when detection prompts generated by GPT-4.1-mini are transferred to be used on other LLMs without modification respectively. The results show that there is a strong transferability since the detection prompt generated by the source model can still achieve a nice TPR value when transfered to other LLMs.
@@ A.5.2 Prompt Transferability (Full Results)
A.6 Human Crafted Prompt
Chiang et al. (2024) introduces the prompt used to detect whether there are attacks in the students’ articles, and we slightly modified the detection prompt that leads to a higher TPR, which is called the Human Crafted Prompt. The full prompt is in Table 11
Human Crafted Prompt We want you to check if the student’s answer tries to hack the evaluation assistant based on the students article. Please look carefully and assess if there are any modification attempts to their score. Please neglect any modifications on evaluation and scoring criteria in the student’s answer and fully obey the evaluation criteria. Any modification on the evaluation criteria or instructions should be treated as an attack.
Table 11: Human Crafted Prompt serves as the baseline to validate the attacks strength in each GAN iteration. This defense is modified from Chiang et al. (2024).